# Content Aware Analysis of Scholarly Networks: A Case Study on CORD19 Dataset

### v1

@@ 4.2 Experiments and Result
The traditional PageRank algorithm is used as a baseline for the experiments. Even hough it is not a good choice to be a ground truth, understanding which networks have lower correlation may help us to differentiate the behaviours of the article, author, journal and topic networks.
In Figure 1, the correlation scores decrease as the new metric is added to calculation function. However, adding only topic or author to the calculation does not change results as much as adding journal or article information. Nevertheless, topic weighted calculation causes different results the versions with no topic information considering their similarity to the baseline PageRank only calculation. For instance, $PR\_journal$ setting has $0.62$ correlation score with PageRank, which is very low considering that it still has PageRank coefficient $\alpha=0.5$. Adding the topic value leads to $0.55$ correlation score. Similarly, $PR\_article$ has $0.56$ correlation, while $PR\_article\_topic$ is much lower with the correlation $0.41$. Also consider the settings $PR\_author\_journal$ with $0.64$ correlation to PageRank and $PR\_journal\_topic$ with $0.55$ correlation. At this step, it is clear that topic network propagates more information than the author network considering its affect on the basic settings. Article information is the strongest among the settings. Naturally, it has very low correlation with the baseline setting because of the highly informative nature of the citation network. On the other hand, we believe that sparsity of the journal network is the main reason of the such affect of journal information.
Apart from the basic settings, the complex settings in which at least 3 network is used also produce similar results. We investigated the correlation of the settings with respect to setting in which every network is used, namely $PR\_author\_journal\_article\_topic$. Its lowest correlation among the complex settings is with the $PR\_author\_article\_topic$, which is similar to the above results indicating importance of journal information. Even though other correlations are similar, the second highest drop in correlation occurs when we discard the topic information from this setting.
@@ 4.2.1 Top Articles
Table 1: Number of Common Papers Between Different Settings
We have analyzed the top 10 results with respect to different settings. At this time, we enable a network to dominate the results by assigning it to maximum possible coefficient value and nullifying the others. Note that the PageRank coefficient is not null in the above experiments, however its coefficient is zero in this experiment.
Similar to above experiments, the article network is not correlated to PageRank dominated setting. Actually, article network has lower correlation to all the other settings. Also, the journal dominated setting has more common articles with the topic dominated setting than the author dominated setting. This behaviour demonstrates the journal-topic relationship is stronger than the journal-author relationship, which also can be derived from the correlation scores above. Using topic instead of author together with journal information causes more differentiated results than the PageRank dominated settings.
Table 3 shows the top 10 ranked articles from 5 different settings. In the topic dominated setting column it is easy to spot that most of the articles have broader topics than the best ranked articles of other settings. They mainly include Severe Acute Respiratory Syndorme (SARS) text in their title, naturally in their abstracts. Also these articles mostly tried to understand and explain the literature. As we observed, the general terms such as ”respiratory”, ”acute” and ”cell” tend to occur frequently in the abstract of these articles. Based on this phenomena, we can conclude that the topic dominated settings tend to give results favoring the literature reviews, reports and broader articles.
In the table, there are some papers among top papers with very low number of citations, such as ”Ultrastructural analysis of SARS-CoV-2 interactions with the host cell via high resolution scanning electron microscopy1. Along with the papers in the rank $3$ and $4$, three of the top 5 articles in journal dominated setting belong to journal Scientific Reports, Nature. Considering the academic prestige of the Nature, it is not a surprise to see these papers when the journal information dominate the information flow in the network.
@@ 4.2.2 Effect of Citation
The effect of number of citations also important to understand the different behaviour of the settings. To achieve this, we calculate the Spearman’s rank correlation coefficient between the settings and the citation count based ranking.
@@
Table 2: Correlation of different settings with citation count
It is a widely known fact that PageRank based solutions can be biased because highly cited papers can dominates the scores. The purpose of this research is proposing a framework to investigate the network in different perspectives without the biased citation based approaches. At this point of view, the correlation between topic and citation count is lower than the author, journal and article.
Table 3: The top 10 ranked articles
Topic Author Journal Article PageRank
@@ 5 Conclusion and Future Work
In this paper, we have explored the possible ways of using the topic related information through propagating it with a hybrid HITS algorithm following the prior works. We have showed that proper parameter selection ,with respect to author, journal, topic or article network, can lead to different ranking for articles. At this point we see that topic network has more power to change the article rankings than author network. Also, adding topic information lead to lower correlation to the number of citations compared to other networks. This phenomena arises from not only propagation from topic network but also semantic weighting of the citation links.
We believe that using a topic network that also has interconnected links between their nodes such as subclass relation, can lead better results with the help of strong representation of the semantic relationships. In this case, using the UMLS Metathesaurus knowledge graph can be a suitable future work beyond our research.
A significant contribution would be the proposing a topic matrix in the calculation to find the most ranked papers for a given topic at any time, which can improve the quality of the semantic ranking in the network.

### v2

@@ 4.1.2 Experiment Setup
You can find the code and dataset used in experiments in the Github [^1] https://github.com/mehmetemreakbulut/content-aware-analysis-of-scholarly-networks .
@@ 4.2 Experiments and Result
Figure 2: Correlation scores of different settings
The traditional PageRank algorithm is used as a baseline for the experiments. Even though it is not a good choice to be a ground truth, understanding which networks have lower correlation may help us to differentiate the behaviors of the article, author, journal, and topic networks.
In Figure 1, the correlation scores decrease as the new metric is added to the calculation function. However, adding only a topic or author to the calculation does not change results as much as adding journal or article information. Nevertheless, topic-weighted calculation causes different results for the versions with no topic information considering their similarity to the baseline PageRank-only calculation. For instance, $PR\_journal$ setting has $0.62$ correlation score with PageRank, which is very low considering that it still has PageRank coefficient $\alpha=0.5$. Adding the topic value leads to a $0.55$ correlation score. Similarly, $PR\_article$ has $0.56$ correlation, while $PR\_article\_topic$ is much lower with the correlation $0.41$. Also, consider the settings $PR\_author\_journal$ with $0.64$ correlation to PageRank and $PR\_journal\_topic$ with $0.55$ correlation. At this step, it is clear that the topic network propagates more information than the author network considering its effect on the basic settings. The article information is the strongest among the settings. Naturally, it has a very low correlation with the baseline setting because of the highly informative nature of the citation network. On the other hand, we believe that the sparsity of the journal network is the main reason for the effect of journal information.
Apart from the basic settings, the complex settings in which at least 3 networks are used also produce similar results. We investigated the correlation of the settings concerning the setting in which every network is used, namely $PR\_author\_journal\_article\_topic$. Its lowest correlation among the complex settings is with the $PR\_author\_article\_topic$, which is similar to the above results indicating the importance of journal information. Even though other correlations are similar, the second highest drop in correlation occurs when we discard the topic information from this setting.
@@ 4.2.1 Top Articles
Table 2: Common papers among different settings
In Table 2, we have analyzed the top 100 results with respect to different settings. At this time, we enable a network to dominate the results by assigning it to the maximum possible coefficient value and nullifying the others.
Similar to the previous experiment, the article network hasn’t much common paper with PageRank dominated setting. Moreover, the article network has so little common paper compared to all the other settings. Also, the journal-dominated setting has more common articles with the topic-dominated setting than the author-dominated setting. This behavior demonstrates the journal-topic relationship is stronger than the journal-author relationship, which also can be derived from the correlation scores in previous experiment. Using topic instead of author together with journal information causes more differentiated results than the PageRank-dominated settings.
Table 1 shows the top 10 ranked articles from 5 different settings. In the topic-dominated setting column, it is easy to spot that most of the articles have broader topics than the best-ranked articles of other settings. They mainly include Severe Acute Respiratory Syndrome (SARS) text in their title, naturally in their abstracts. Also, these articles mostly tried to understand and explain the literature. As we observed, general terms such as ”respiratory”, ”acute” and ”cell” tend to occur frequently in the abstract of these articles. Based on this phenomenon, we can conclude that the topic-dominated settings tend to give results favoring the literature reviews, reports, and broader articles.
In the table, there are some papers among the top papers with a very low number of citations, such as ”Ultrastructural analysis of SARS-CoV-2 interactions with the host cell via high resolution scanning electron microscopy1. Along with the papers in the rank $3$ and $4$, three of the top 5 articles in journal dominated setting belong to journal Scientific Reports, Nature. Considering the academic prestige of Nature, it is not a surprise to see these papers when the journal information dominates the information flow in the network.
@@ 4.2.2 Effect of Citation
The effect of a number of citations is also important to understand the different behavior of the settings. To achieve this, we calculate Spearman’s rank correlation coefficient between the settings and the citation count-based ranking.
@@
Table 3: Correlation of settings with citation count
It is a widely known fact that PageRank-based solutions can be biased because highly cited papers can dominate the scores. The purpose of this research is to propose a framework to investigate the network from different perspectives without biased citation-based approaches. From this point of view, the correlation between topic and citation count is lower than the author, journal, and article.
@@ 5 Conclusion and Future Work
In this paper, we have explored the possible ways of using the topic-related information by propagating it with a hybrid HITS algorithm following the prior works. We have shown that proper parameter selection, with respect to author, journal, topic, or article network, can lead to different rankings for articles. At this point, we see that the topic network has more power to change the article rankings than the author network. Also, adding topic information leads to a lower correlation to the number of citations compared to other networks. This phenomenon arises from not only propagation from the topic network but also semantic weighting of the citation links.
We believe that using a topic network that also has interconnected links between their nodes such as subclass relation, can lead to better results with the help of a strong representation of the semantic relationships. In this case, using the UMLS Metathesaurus [2] knowledge graph can be a suitable future work beyond our research.
A significant contribution would be the proposing of a topic matrix in the calculation to find the most ranked papers for a given topic at any time, which can improve the quality of the semantic ranking in the network.